# Revisiting Jahn–Teller Transitions in Correlated Oxides with Monte Carlo Modeling

### v2

@@ II.1 Hamiltonian
We define a global classical, non-quantum Hamiltonian, for a system with N JT-active octahedra, as follows:
@@ II.1.1 Single-ion term
in which $\rho_{i}$ is the magnitude of the JT distortion at the ith JT-active site. $\alpha$ and $\beta$ are empirical parameters which determine the functional dependence of enthalpy on $\rho$. Typically, we would expect $\alpha<0$, $\beta>0$, and $|\beta|<|\alpha|$.
Figure 3: The results of a Monte Carlo refinement with dynamic $\rho$ ($p_{\mathrm{switch}}=1/2$), reproducing Phase 5 (C-type) of the Ahmed and Gehring Ahmed and Gehring (2005) phase diagram for the anisotropic Potts model in a $8\times 8\times 8$ perovskite supercell. (a) Example configuration in the $bc$-plane in a randomly-selected cross-section. (b) Example configuration in the $ac$-plane in a randomly-selected cross-section. Energy parameters $\alpha=-1$$\beta=1/2$ from equation 1. Simulated annealing was used before settling on a final $T=0.001$. This run lasted for $10^{7}$ iterations. Plots of energy and $\langle\rho\rangle$ with iteration can be seen in Figure S12.
@@ II.1.2 Geometry term
where $K_{\mathrm{oxy}}$ is an empirical parameter, $K_{\mathrm{oxy}}>0$, which determines the strength of this O under-bonding penalty, with the implication of this equation shown in Figure 2(b). The “underbonding density” term $s_{O}$ is given by summing over all three nearest-neighbour Ni cations to a given O anion thus:
@@
We note that the nature of the Hamiltonian we use here renders us insensitive to the possibility of disproportionation as proposed in some prior works Chen et al. (2011a); Foyevtsova et al. (2019); Poletayev et al. (2025).
@@ II.2 Model cell
Simulations are run over periodic cells in two- or three-dimensions for the layered nickelate and perovskite case, respectively. For the perovskite case, a cubic primitive cell is used with a $8\times 8\times 8$ grid of metal cations (512 atoms). For the layered nickelate case, the cell is a $10\times 30$ rhombus-shaped grid of Ni cations (300 atoms), where the two basis vectors enclose 60. The use of a two-dimensional lattice relies on the axiom that inter-layer interactions are not significant for the JT ordering. Unless otherwise indicated, the starting configuration for all runs was one in which JT distortions were randomly distributed within the unit cell, all with identical magnitude $\rho=1$.
@@ II.3 Metropolis algorithm
The Monte Carlo simulation runs over a number of iterations ($10^{5}$ to $10^{7}$) using the Metropolis algorithm Metropolis et al. (1953). At each iteration, a metal site is selected at random. With probability $0\leq p_{\mathrm{switch}}\leq 1$, a $\rho$ shift occurs in which the magnitude of $\rho$ is varied randomly by some amount, with $0.1\leq\Delta\rho\leq 0.1$ (with $\rho$ capped at 0 at the lower limit and 1.5 at the upper limit). Alternatively, a direction shift occurs, in which the axis of elongation of the metal site is shifted randomly. Energy is then re-calculated according to Equation 1. If $\Delta E\leq 0$, the change is accepted. Otherwise, the change is accepted with probability $\exp\left[-\Delta E/T\right]$. For most testing in this work, $p_{\mathrm{switch}}$ was set arbitrarily at 0.5.
@@ III Results: JT-distorted perovskite
Figure 4: The mean energy terms $E_{\mathrm{total}}$, $E_{\mathrm{Geometric}}$ (Eq 3), and $E_{\mathrm{single-ion}}$ (Eq 2), heat capacity $C$, and mean $\langle\rho\rangle$, averaged over the final 10% of iterations in Monte Carlo simulations, as a function of temperature for the collinear and C-type orbital orderings in a $8\times 8\times 8$ perovskite lattice. The Monte Carlo simulations ran for $10^{7}$ iterations in total at each temperature. This figure is reproduced on a non-logarithmic temperature scale in SI Figure S13.
Before allowing $\rho$ to vary (i.e. by setting all metal sites to have constant non-zero JT magnitude, and setting $p_{\mathrm{switch}}=0$), we tested our model where the Hamiltonian consisted of the $E_{\mathrm{geometry}}^{\mathrm{perovskite}}$ term only (i.e. $\alpha=\beta=0$), as shown in SI Section S2. This reproduced the $(J_{1},J_{2})$ anisotropic Potts model phase diagram of Ahmed and Gehring Ahmed and Gehring (2005), validating their results and our simulation approach.
@@
Figure 6: An example converged configuration of a $10\times 15$ layered nickelate. This simulation used a static $\rho$ ($p_{\mathrm{\rho}}=0$; $\alpha=\beta=0$) model, with $K_{\mathrm{oxy}}=10^{10}$ and $T=0.001$ after simulated annealing (with $T_{\mathrm{max}}=10$), and was run for $10^{6}$ iterations. Collinear domains can be seen, often bound by defects of over-bonded (no JT axes) or under-bonded (multiple JT axes) oxygen anions, indicating that this is a local, but not global, energy minimum. In this simulation, we set $K_{\mathrm{Oxy}}=10^{10}$.
@@
We next consider the variable-temperature behaviour of these orbital orderings. We performed simulations, with simulated annealing, at a series of temperatures up to $T=30$, for both the collinear and C-type ($J_{1}$$J_{2}$) configurations. Figure 4 shows the mean energy and $\langle\rho\rangle$ for the final 10% of iterations in the simulations as a function of temperature, for both cases. We see that, as temperature increases, there is a transition, which can be seen best in the heat capacity data [Figure 4], where heat capacity is the variance in $E_{\mathrm{total}}$ divided by $T^{2}$. $\langle\rho\rangle$ decreases with heating, with the rate of decrease being low on approach to the transition, and then a rapid decrease before settling at around $\langle\rho\rangle\approx 0.8$. We have tested the resilience of this finding against varying cell size [Figures S16 and S17] with little variation. The finding that $\langle\rho\rangle$ decreases on heating even in the low-temperature regime approximately corresponds to experimental observations from diffraction on LaMnO3, although KCuF3 exhibits the opposite trend Marshall et al. (2013).
A convergence of $\langle\rho\rangle\rightarrow 0.8$ at high temperatures appears consistent with the experimental reports (from PDF and EXAFS) of JT-like octahedral distortion at high-temperatures, if we interpret it as the system remaining in the enthalpic minimum imposed by Eq 2. However, when we plot histograms of $\rho$ in the final configuration [Figure 5], we see that the distribution at high temperature does not resemble the distribution at low temperature, which would be the signature of an order-disorder transition. Instead, we see $N(\rho)$ appears invariant with the magnitude, which suggests that rather than a persistence of JT distortions the system is merely dynamically exploring the available phase space of possible values of $\rho$, which is consistent with the recent theoretical study of LaMnO3 Batnaran et al. (2025). These findings could be interpreted more in terms of a displacive transition than an order-disorder transition, although the invariance in $N(\rho)$ with $\rho$ indicates that if there is a free energy minimum at the origin in $E_{g}(Q_{2},Q_{3})$ phase space, it is not significantly more stable than other regions of the phase space.
@@ IV Results: Layered nickelate
We next simulated the layered nickelates. Here, we rely on the axiom that inter-layer effects are negligible due to the AO6 octahedral layer (AA=Li,Na), and consider only single layers. As with the perovskites, we first performed Monte Carlo simulations for the static $\rho$ case where all sites begin JT-distorted, with $p_{\mathrm{switch}}=0$, and the Hamiltonian consisting solely of the $E_{\mathrm{geometry}}^{\mathrm{nickelate}}$ term (i.e. $\alpha=\beta=0$), with $T<<1$. An example of a converged layer configuration is shown in Figure 6. We see the emergence of ordered collinear domains, as reported from diffraction studies on NaNiO2 Dyer et al. (1954); Dick et al. (1997); Sofin and Jansen (2005); Nagle-Cocco et al. (2022) and stoichiometric LiNiO2 Phillips et al. (2025). The boundary between these domains resembles the “zigzag” ordering which has been proposed as a possible ground state for LiNiO2 on the basis of DFT studies Chen et al. (2011b); Foyevtsova et al. (2019). Defects then occur when two or three domains intersect. Such defects involve either over-bonded (no JT axes) or under-bonded (multiple JT axes) oxygen anions, and resemble the dimer or even trimer models proposed by Chung et al (2005) Chung et al. (2005) [Figure S1] which have been found from DFT to be highly unstable Chen et al. (2011b). This domain effect, in contrast, does not seem to be important for the perovskites, likely because they can relieve frustration by stacking patterns along $c$, which reduces the likelihood of frustrated interactions.
@@ V Conclusion
In this study, we successfully extended the Monte Carlo model of Ahmed and Gehring Ahmed and Gehring (2005, 2006, 2009) to include variable Jahn-Teller (JT) distortion amplitudes, enabling us to retain the simplicity of the Ahmed and Gehring model, while allowing for the possibility that JT distortions may vanish locally at high-temperature. The results of this approach are qualitatively consistent with both experimental evidence for persistence of JT-like distortion modes above JT transitions in perovskites Qiu et al. (2005); Chatterji et al. (2003) (which, in our simulations, manifests as a high $\langle\rho\rangle$), and recent computational results which suggest that the high-temperature state of JT-distorted perovskites is not characterised by well-defined JT-distorted octahedra Batnaran et al. (2025), but rather a dynamic sampling of the ($Q_{2}$$Q_{3}$) phase space. These findings suggest that a displacive-like transition, such as that recently reported experimentally for the cJT suppression in the layered nickelates Genreith-Schriever et al. (2024); Nagle-Cocco et al. (2024), may in fact be a more accurate description of cJT suppression than the more orthodox order-disorder description.
Despite our interpretation that both the transition in perovskites and layered nickelates may be better described as displacive than order-disorder, there are clearly differences between the two systems due to the geometry of the lattice. We reported that the converged $\langle\rho\rangle$ of the perovskite lattices is much higher ($\sim$0.8) than the nickelate lattice ($\sim$0.4) at high temperature, and as shown in Figure 5 the number distribution of distortion magnitudes appears invariant with $\rho$ for the perovskite case but levels off as $\rho$ increases for the nickelate case. This is likely a consequence of the subextensivity of configurational entropy of disordered octahedral elongations on the nickelate lattice, as previously reported Nagle-Cocco et al. (2024). The differences likely manifest experimentally in a larger peak skewness, for the first metal-oxygen peak in experimental pair distribution function data, for perovskites like LaMnO3 compared with nickelates like NaNiO2. This may be a factor in the historic attribution of order-disorder behaviour to the JT transitions in the perovskites.
@@
Code was written in Python 3 Van Rossum (1995), with assistance from generative large-language models ChatGPT OpenAI et al. (2024) version 5 and GitHub Copilot. Graphs are prepared using Matplotlib Hunter (2007). Crystal structures prepared using VESTA 3 Momma and Izumi (2011).
@@ Appendix A Simulated annealing
Figure S2: The orbital ordering in the perovskite La0.5Ba0.5CoO3, as reported for (a) ordered Nakajima et al. (2005) and (b) disordered Fauth et al. (2002) A-site cations. Here, A-site cations are excluded, blue cations are Co, grey bonds are short and pink bonds are long.
@@ 4 $J_{2}>0,\;\;J_{1}>0$ $J_{1}=J_{2}=1$ 15 Figure S8 -
Beyond this, we then tested a dynamic $\rho$ model, in which $\alpha\neq 0$ and $\beta\neq 0$, and $p_{\mathrm{switch}}=0.5$, to ensure our extended Hamiltonian in Equation 1 can also reproduce the phase diagram. We test this on phases 1 and 5, shown in Figures S11 and 3 respectively.
@@
Figure S13: The mean energy terms $E_{\mathrm{total}}$, $E_{\mathrm{Geometric}}$ (Eq 3), and $E_{\mathrm{single-ion}}$ (Eq 2), heat capacity $C$, and $\langle\rho\rangle$, averaged over the final 10% of iterations in Monte Carlo simulations, as a function of temperature for the collinear and C-type orbital orderings in a $8\times 8\times 8$ perovskite lattice. The Monte Carlo simulations ran for $10^{7}$ iterations in total at each temperature. This figure is reproduced on a logarithmic temperature scale in Figure 4.
@@
To test the impact of supercell size on the results (and ensure the resilience of our findings using the $8\times 8\times 8$ supercell with the perovskite structure) we also performed measurements using supercells with edge size 4 and 6. Figures S16 and S17 show the variable-temperature energy, heat capacity, and $\langle\rho\rangle$ for $4\times 4\times 4$ and $6\times 6\times 6$ supercells, respectively.

### v3

@@ II.1 Hamiltonian
We define a global classical Hamiltonian, for a system with N JT-active octahedra, as follows:
@@ II.1.1 Single-ion term
in which $\rho_{i}$ is the magnitude of the JT distortion at the ith JT-active site. $\alpha$ and $\beta$ are empirical parameters which determine the functional dependence of enthalpy on $\rho$. Typically, we would expect $\alpha<0$, $\beta>0$, and $|\beta|<|\alpha|$. In the simulations presented here, where variable-magnitude $\rho$ is enabled, $\alpha=1$ and $\beta=-1/2$.
Figure 3: The results of a Monte Carlo refinement with dynamic $\rho$ ($P_{\mathrm{switch}}=1/2$), reproducing Phase 5 (C-type) of the Ahmed and Gehring Ahmed and Gehring (2005) phase diagram for the anisotropic Potts model in an $8\times 8\times 8$ perovskite supercell. (a) Example configuration in the $bc$-plane in a randomly-selected cross-section. (b) Example configuration in the $ac$-plane in a randomly-selected cross-section. Energy parameters $\alpha=-1$$\beta=1/2$ from equation 1. Simulated annealing was used before settling on a final $T=0.001$. This run lasted for $10^{7}$ iterations. Plots of energy and $\langle\rho\rangle$ with iteration can be seen in Figure S12.
@@ II.1.2 Geometry term
where $K_{\mathrm{oxy}}$ is an empirical parameter, $K_{\mathrm{oxy}}>0$, which determines the strength of this O under-bonding penalty, with the implication of this equation shown in Figure 2(b) assuming all $\rho$ are 1. The “underbonding density” term $s_{O}$ is given by summing over all three nearest-neighbour Ni cations to a given O anion thus:
@@
We note that the nature of the Hamiltonian we use here renders us insensitive to the possibility of disproportionation as proposed in some prior works for both layered nickelates Chen et al. (2011a); Foyevtsova et al. (2019); Poletayev et al. (2025) and perovskites García-Muñoz et al. (1994); Mizokawa et al. (2000); Garcia-Munoz et al. (2009); Pascut and Haule (2023).
@@ II.2 Model cell
Simulations are run over periodic cells in two- or three-dimensions for the layered nickelate and perovskite case, respectively. For the perovskite case, a cubic primitive cell is used with an $8\times 8\times 8$ grid of metal cations (512 atoms). For the layered nickelate case, the cell is a $10\times 30$ rhombus-shaped grid of Ni cations (300 atoms), where the two basis vectors enclose 60. The use of a two-dimensional lattice for the layered nickelates relies on the axiom that inter-layer interactions are not significant for the JT ordering. Unless otherwise indicated, the starting configuration for all runs was one in which JT distortions were randomly distributed within the unit cell, all with identical magnitude $\rho=1$.
@@ II.3 Metropolis algorithm
The Monte Carlo simulation runs over a number of iterations ($10^{5}$ to $10^{7}$) using the Metropolis algorithm Metropolis et al. (1953). At each iteration, a metal site is selected at random. With probability $0\leq P_{\mathrm{switch}}\leq 1$, a $\rho$ shift occurs in which the magnitude of $\rho$ is varied randomly by some amount, with $0.1\leq\Delta\rho\leq 0.1$ (with $\rho$ capped at 0 at the lower limit and 1.5 at the upper limit). Alternatively, with probability $P_{\mathrm{flip}}=1-P_{\mathrm{switch}}$, a direction shift occurs, in which the axis of elongation of the metal site is shifted randomly. Either way, energy is then re-calculated according to Equation 1. If $\Delta E\leq 0$, the change is accepted. Otherwise, the change is accepted with probability $\exp\left[-\Delta E/T\right]$. For most testing in this work, $P_{\mathrm{switch}}$ was set arbitrarily at 0.5.
@@ III Results: JT-distorted perovskite
Figure 4: The mean energy terms $E_{\mathrm{total}}$, $E_{\mathrm{Geometric}}$ (Eq 3), and $E_{\mathrm{single-ion}}$ (Eq 2), heat capacity $C$, and mean $\langle\rho\rangle$, averaged over the final 10% of iterations in Monte Carlo simulations, as a function of temperature for the collinear and C-type orbital orderings in an $8\times 8\times 8$ perovskite lattice. The Monte Carlo simulations ran for $10^{7}$ iterations in total at each temperature. This figure is reproduced on a non-logarithmic temperature scale in SI Figure S13.
Before allowing $\rho$ to vary (i.e. by setting all metal sites to have constant non-zero JT magnitude, and setting $P_{\mathrm{switch}}=0$), we tested our model where the Hamiltonian consisted of the $E_{\mathrm{geometry}}^{\mathrm{perovskite}}$ term only (i.e. $\alpha=\beta=0$), as shown in SI Section S2. This reproduced all phases of the $(J_{1},J_{2})$ anisotropic Potts model phase diagram of Ahmed and Gehring Ahmed and Gehring (2005), validating their results and our simulation approach.
@@
Figure 6: An example converged configuration of a $10\times 15$ layered nickelate. This simulation used a static $\rho$ ($P_{\mathrm{switch}}=0$; $\alpha=\beta=0$) model, with $K_{\mathrm{oxy}}=10^{10}$ and $T=0.001$ after simulated annealing (with $T_{\mathrm{max}}=10$), and was run for $10^{6}$ iterations. Collinear domains can be seen, often bound by defects of over-bonded (no JT axes) or under-bonded (multiple JT axes) oxygen anions, indicating that this is a local, but not global, energy minimum. In this simulation, we set $K_{\mathrm{Oxy}}=10^{10}$.
@@
We next consider the variable-temperature behaviour of these orbital orderings. We performed simulations, with simulated annealing, at a series of temperatures up to $T=30$, for both the collinear and C-type ($J_{1}$$J_{2}$) configurations. Figure 4 shows the mean energy and $\langle\rho\rangle$ for the final 10% of iterations in the simulations as a function of temperature, for both cases. We see that, as temperature increases, there is a transition, which can be seen best in the heat capacity data [Figure 4], where heat capacity is the variance in $E_{\mathrm{total}}$ divided by $T^{2}$. $\langle\rho\rangle$ decreases with heating, with the rate of decrease being low on approach to the transition, and then a rapid decrease before settling at around $\langle\rho\rangle\approx 0.8$. We have tested the resilience of this finding against varying cell size [Figures S16 and S17] with little variation. The finding that $\langle\rho\rangle$ decreases on heating even in the low-temperature regime approximately corresponds to experimental observations from diffraction on LaMnO3 that $d\rho/dT<0$, although KCuF3 exhibits the opposite trend Marshall et al. (2013).
A convergence of $\langle\rho\rangle\rightarrow 0.8$ at high temperatures appears consistent with the experimental reports (from PDF and EXAFS) of JT-like octahedral distortion at high temperatures, if we interpret it as the system remaining in the enthalpic minimum imposed by Eq 2. However, when we plot histograms of $\rho$ in the final configuration [Figure 5], we see that the distribution at high temperature does not resemble the distribution at low temperature, which would be the signature of an order-disorder transition. Instead, we see $N(\rho)$ appears invariant with the magnitude, which suggests that rather than a persistence of JT distortions the system is merely dynamically exploring the available phase space of possible values of $\rho$, which is consistent with the recent theoretical study of LaMnO3 Batnaran et al. (2025). These findings could be interpreted more in terms of a displacive transition than an order-disorder transition, although the invariance in $N(\rho)$ with $\rho$ indicates that if there is a free energy minimum at the origin in $E_{g}(Q_{2},Q_{3})$ phase space, it is not significantly more stable than other regions of the phase space.
@@ IV Results: Layered nickelate
We next simulated the layered nickelates. Here, we rely on the axiom that inter-layer effects are negligible due to the alternating AO6 octahedral layers (AA=Li,Na), and consider only single layers. As with the perovskites, we first performed Monte Carlo simulations for the static $\rho$ case where all sites begin JT-distorted, with $P_{\mathrm{switch}}=0$, and the Hamiltonian consisting solely of the $E_{\mathrm{geometry}}^{\mathrm{nickelate}}$ term (i.e. $\alpha=\beta=0$), with $T<<1$. An example of a converged layer configuration is shown in Figure 6. We see the emergence of ordered collinear domains, as reported from diffraction studies on NaNiO2 Dyer et al. (1954); Dick et al. (1997); Sofin and Jansen (2005); Nagle-Cocco et al. (2022) and stoichiometric LiNiO2 Phillips et al. (2025). The boundary between these domains resembles the “zigzag” ordering which has been proposed as a possible ground state for LiNiO2 on the basis of DFT studies Chen et al. (2011b); Foyevtsova et al. (2019). Defects then occur when two or three domains intersect. Such defects involve either over-bonded (no JT axes) or under-bonded (multiple JT axes) oxygen anions, and resemble the dimer or even trimer models proposed by Chung et al (2005) Chung et al. (2005) [Figure S3(e,h)] which have been found from DFT to be highly unstable Chen et al. (2011b). This domain effect, in contrast, does not seem to be important for the perovskites, likely because they can relieve frustration by stacking patterns along $c$, which reduces the likelihood of frustrated interactions.
@@ V Conclusion
In this study, we successfully extended the Monte Carlo model of Ahmed and Gehring Ahmed and Gehring (2005, 2006, 2009) to include variable Jahn–Teller (JT) distortion amplitudes, enabling us to retain the simplicity of the Ahmed and Gehring model, while allowing for the possibility that JT distortions may vanish locally at high temperature. The results of this approach are qualitatively consistent with both experimental evidence for persistence of JT-like distortion modes above JT transitions in perovskites Qiu et al. (2005); Chatterji et al. (2003) (which, in our simulations, manifests as a high $\langle\rho\rangle$), and recent computational results which suggest that the high-temperature state of JT-distorted perovskites is not characterised by well-defined JT-distorted octahedra Batnaran et al. (2025), but rather a dynamic sampling of the ($Q_{2}$$Q_{3}$) phase space. These findings suggest that a displacive-like transition, such as that recently reported experimentally for the cJT suppression in the layered nickelates Genreith-Schriever et al. (2024); Nagle-Cocco et al. (2024), may in fact be a more accurate description of cJT suppression than the more orthodox order-disorder description.
Despite our interpretation that both the transition in perovskites and layered nickelates may be better described as displacive than order-disorder, there are clearly differences between the two systems due to the geometry of the lattice. We reported that the converged $\langle\rho\rangle$ of the perovskite lattices is much higher ($\sim$0.8) than the nickelate lattice ($\sim$0.4) at high temperature, and as shown in Figure 5 the number distribution of distortion magnitudes appears invariant with $\rho$ for the perovskite case but decreases as $\rho$ increases for the nickelate case. This is likely a consequence of the subextensivity of configurational entropy of disordered octahedral elongations on the nickelate lattice, as previously reported Nagle-Cocco et al. (2024). The differences likely manifest experimentally in a larger skewness for the first metal-oxygen peak in the experimental pair distribution function for perovskites like LaMnO3, as compared with nickelates like NaNiO2. This may be a factor in the historic attribution of order-disorder behaviour to the JT transitions in the perovskites.
@@
Code was written in Python 3 Van Rossum (1995), with assistance from generative large-language models ChatGPT OpenAI et al. (2024) version 5. Graphs are prepared using Matplotlib Hunter (2007). Crystal structures prepared using VESTA 3 Momma and Izumi (2011).
@@ Appendix A Simulated annealing
In this section, some other orbital orderings reported in the literature are presented, to support the discussion throughout the introduction and results of the main text.
@@
Figure S2: The orbital ordering of the JT distortion in the perovskite La0.5Ba0.5CoO3, as reported for (a) ordered Nakajima et al. (2005) and (b) disordered Fauth et al. (2002) A-site cations. In both schematics, the crystal structure is viewed along the [111] direction. Here, A-site cations are excluded. Blue atoms represent Co cations and red atoms reperesent O anions. In terms of bonds, grey bonds are short and pink bonds are long.
@@ 4 $J_{2}>0,\;\;J_{1}>0$ $J_{1}=J_{2}=1$ 15 Figure S8 -
Beyond this, we then tested a dynamic $\rho$ model, in which $\alpha\neq 0$ and $\beta\neq 0$, and $P_{\mathrm{switch}}=0.5$, to ensure our extended Hamiltonian in Equation 1 can also reproduce the phase diagram. We test this on phases 1 [Figure S11] and 5 [Figures 3 and S12].
@@
In the main text, we presented plots of energy, heat capacity, and mean JT magnitude $\langle\rho\rangle$ as a function of temperature, with temperature presented on a logarithmic scale. In this section, we present the same figures but on a non-logarithmic scale for reference. Figure S13 shows the results for collinear and C-type orbital ordering perovskites, Figure S14 shows the case for nickelates with $K_{\mathrm{oxy}}=1,10,10^{10}$, and Figure S15 shows the combined case.
Figure S13: The mean energy terms $E_{\mathrm{total}}$, $E_{\mathrm{Geometric}}$ (Eq 3), and $E_{\mathrm{single-ion}}$ (Eq 2), heat capacity $C$, and $\langle\rho\rangle$, averaged over the final 10% of iterations in Monte Carlo simulations, as a function of temperature for the collinear and C-type orbital orderings in an $8\times 8\times 8$ perovskite lattice. The Monte Carlo simulations ran for $10^{7}$ iterations in total at each temperature. This figure is reproduced on a logarithmic temperature scale in Figure 4.
@@
To test the impact of supercell size on the results (and ensure the resilience of our findings using the $8\times 8\times 8$ supercell with the perovskite structure) we also performed measurements using supercells with edge size 4 and 6. Figures S16 and S17 show the variable-temperature energy, heat capacity, and $\langle\rho\rangle$ for $4\times 4\times 4$ and $6\times 6\times 6$ supercells, respectively. We did not observe qualitative variation in results with cell size.
@@
In Figure 5 in the main text, we presented histograms of $\rho$ in the final configuration for a $8\times 8\times 8$ collinear perovskite ordered supercell and a $10\times 30$ nickelate layer with $K_{\mathrm{oxy}}\rightarrow\infty$. In this section, we present such histograms for both perovskite orderings studied with temperature, along with all three values of $K_{\mathrm{oxy}}$ used in this work.